# Decomposing Centrally Symmetric Convex Polyhedral Surfaces into Parallelograms

### 22 Mar 2026 v1

@@ Abstract
Let $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the moduli space of centrally symmetric convex polyhedral surfaces with $2N$ labeled vertices and prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$. We show that $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ has the structure of a real hyperbolic manifold of dimension 2N32N-3. When $N=4$ and $5$, we show that every surface in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ can be decomposed into at most $2\binom{2N-2}{2}$ parallelograms, and the decomposition is invariant under the antipodal map. Using the edge-lengths of these parallelograms as coordinates, we show that the moduli space of centrally symmetric polyhedral surfaces with $8$ unlabeled vertices and cone-deficits $\frac{\pi}{2}$ is isometric to the quotient of a real hyperbolic regular ideal $5$-simplex by the dihedral group $D_{6}$.
@@ 1 Introducton
A polyhedron is convex if $0<\delta_{i}<2\pi$ for all $i$. In [10], Thurston considered the moduli space of convex polyhedra with prescribed cone-deficits. By decomposing the surfaces of polyhedra into triangles, he built coordinates to describe these polyhedra, and showed that their moduli space admits a Hermitian metric, with respect to which it is a complex hyperbolic manifold. In addition, under certain conditions, the metric completion of this moduli space is an orbifold. This result provides a geometric way to interpret the monodromy groups of hypergeometric functions investigated by Picard [7][8], Mostow and Deligne [2][3][6]. In addition, this result has been applied in counting and enumerating specific structures on the sphere, such as counting convex tilings [5] and enumerating fullerenes [4]. Thurston’s idea in building coordinates and the Hermitian form is also used to study the moduli space of convex polygons [1].
@@
Inspired by this work, we study the space of centrally symmetric convex polyhedra. A polyhedron is centrally symmetric if it is isometric to a polyhedron symmetric about the origin, which means that a point $x$ lies on its surface if and only if its antipodal point x-x does. The map that sends $x$ to x-x for all $x$ on the surface is called the antipodal map. Thus, a centrally symmetric polyhedron has an even number $2N$ of vertices, and its cone-deficits are described by $N$ positive numbers that sum up to $2\pi$.
@@
The main results of this paper can be summarized as follows. In Section 2, we show that the moduli space of centrally symmetric convex polyhedra with prescribed cone-deficits has the structure of a real hyperbolic manifold. The way we compute the signature of the quadratic form is very similar to the way Thurston computed it for the Hermitian form in [10]. In addition, it is known that every centrally symmetric convex polygon can be decomposed into finitely many parallelograms. This leads us to conjecture that a similar result holds for every centrally symmetric convex polyhedral surface. In a previous work [12], we showed that the surface of every centrally symmetric octahedron has a parallelogram decomposition that is invariant under the antipodal map. In Section 3, we establish a method that relates a parallelogram decomposition of a polyhedral surface to a structure called “loop arrangement” on the sphere. Using this method, we generalize this result to centrally symmetric convex polyhedra with $8$ and $10$ vertices in Section 4 and 5. In these cases, it turns out that the edge-lengths of the parallelograms in the decomposition can be used as coordinates to describe polyhedra. These coordinates are different from Thurston’s coordinates in [10], but may sometimes provide a more concrete geometric description of the moduli spaces when the polyhedra are centrally symmetric. For example, in Section 4, we show that the moduli space of centrally symmetric polyhedra with $8$ vertices whose cone-deficits are $\frac{\pi}{2}$ is isometric to the quotient of a real hyperbolic regular ideal $5$-simplex by a group isomorphic to $D_{6}$.
@@ 2 Space of Centrally Symmetric Convex Polyhedra
Let $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$ be $N$ positive numbers that sum up to $2\pi$. Let $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the space of centrally symmetric convex polyhedral surfaces with $N$ pairs of antipodal vertices labeled by $i^{+}$ and $i^{-}$ (where $1\leq i\leq N$) with prescribed cone-deficit $\delta_{i}$. In this space, two surfaces are considered equivalent if and only if they differ by a Euclidean isometry that preserves the labels of all the vertices.
@@
Next, let $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the moduli space of $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$, in which two polyhedral surfaces are equivalent if and only if they differ by a similarity preserving the vertex-labels. Alternatively, we can regard $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ as the space of polyhedral surfaces in $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ whose surface areas are equal to $1$.
In [10], it was shown that the moduli space of convex polyhedra with $N$ prescribed cone-deficits has the structure of a complex hyperbolic manifold of dimension N3N-3. The main idea of proof is to show that the surface area is a Hermitian form of signature (1,n3)(1,n-3) with respect to the coordinates arising from decomposing the surfaces into triangles. The argument is based on induction on the number of vertices, where multiple vertices of a surface may “collide” to produce a surface with fewer vertices. One may refer to [10] and [9] for more details of this operation.
@@
Therefore, surfaces near $\mathscr{S}$ in $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ with area $1$ correspond to vectors in $\mathbb{R}^{2N-2}$ of length $1$ with respect to a quadratic form of signature (1,2N3)(1,2N-3). This establishes local coordinate charts from $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ to the real hyperbolic space $\mathbb{H}^{2N-3}$. Finally, any change-of-coordinate map on $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ keeps the surface area invariant. This shows that $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ is a real hyperbolic manifold of dimension 2N32N-3. ∎
Let $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the metric completion of $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ with respect to the real hyperbolic metric above. For example, the surface in the third picture of Figure 4 is not in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ but in $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ if it has area $1$. This is because one can construct a Cauchy sequence of surfaces in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ converging to it by colliding two vertices and normalizing the surface area at the same time. We will return to this space in Section 4.
@@ 3 Loop Arrangements
Suppose that we are given $2N$ distinct vertices on $S^{2}$ labeled by $^{+}, 11^{-}$, $^{+}, 22^{-}$, $\dots$, $N^{+}$ and $N^{-}$, where $i^{+}$ and $i^{-}$ are antipodal for $1\leq i\leq N$. By a loop we mean a great circle on $S^{2}$ that does not pass through any labeled vertices. A set of loops form a loop arrangement if any two of them are non-homotopic with respect to the labeled vertices. Since loops are great circles, every two of them intersect exactly twice at two antipodal points. We assume without loss of generality that no three loops are concurrent, otherwise we can perturb any one of the them. As an example, in Figure 5, we sketch the “front view” of a loop arrangement with $6$ loops on a sphere with $8$ labeled vertices. Each red curve represents a loop. We use black dots to represent the vertices $1^{+}$, $2^{+}$, $3^{+}$, and $4^{+}$. In addition, we use gray dots to represent the vertices $2^{-}$, $3^{-}$, and $4^{-}$ (even though they should be invisible from the actual front view) because it will be helpful for our demonstrations later.
Figure 5: A loop arrangement with $6$ loops (red) on a sphere with $8$ labeled vertices.
@@
We first label all loops in $\mathscr{A}$. For example, if $\mathscr{A}$ is the loop arrangement in Figure 5, we can denote its $8$ loops by $l_{a}$, $l_{b}$, $l_{c}$, $l_{d}$, $l_{e}$, and $l_{f}$. For convenience, these loops are marked by $a$, $b$, $\dots$, $f$ in the first picture of Figure 6. In the second picture, we draw the dual graph $\mathscr{D}^{*}$ of $\mathscr{A}$ in blue.
@@
Next, we assign angles to each quadrilateral. Let $Q$ be a quadrilateral in $\mathscr{D}^{*}$. Consider two edges of $Q$ that meet at a vertex $v$ in $Q$. We are going to assign a value $\theta$ as the angle made by these two edges in $Q$. To compute $\theta$, let $l_{i}$ and $l_{j}$ be the two loops to which these two edges are transverse. Then $l_{i}$ and $l_{j}$ divide the sphere into $4$ lunes, one of which contains $v$ in its interior. We denote this lune by $\mathcal{L}$. For example, in the first picture of Figure 7, we sketch a quadrilateral $Q$ (blue), the angle $\theta$ we want to compute in $Q$, and the lune $\mathcal{L}$ (red) we consider to compute $\theta$.
@@
Since $2N-2\geq 4$, we can choose three distinct loops in $\mathscr{A}$, denoted by $l_{i}$, $l_{j}$ and $l_{k}$. By assumption, they are not concurrent, so they divide the sphere into $8$ geodesic triangles. Denote one of these geodesic triangles by $\mathcal{T}$. Let $u_{i}$, $u_{j}$, and $u_{k}$ be the vertices of $\mathcal{T}$ such that $u_{i}$ is opposite to the edge of $\mathcal{T}$ contained in $l_{i}$, and similarly for $u_{j}$ and $u_{k}$. The first picture of Figure 8 is a sketch of $\mathcal{T}$ with its vertices.
@@
Now we focus on all the loops in $\mathscr{A}$. These loops divide $S^{2}$ into convex geodesic polygons. Let $\mathcal{P}_{m}$ be one of these polygons. Without loss of generality, we assume that it is bounded by the loops $l_{1}$, $l_{2}$, $\dots$, and $l_{m}$ in $\mathscr{A}$. Figure 9 shows an example when $m=6$. Consider the $m$ parallelograms in $\mathscr{D}^{*}$ dual to the $m$ vertices of $\mathcal{P}_{m}$. Let $v^{*}_{1}$, $v^{*}_{2}$, $\dots$, $v^{*}_{m}$ be the vertices of these parallelograms inside $\mathcal{P}_{m}$, which may be distinct or repeat. Let $\Theta_{m}$ be the sum of the $m$ angles of these parallelograms at the $v^{*}_{i}$’s. In the example in Figure 9, the $6$ parallelograms are drawn in blue, and the sum of all the angles marked in red equals $\Theta_{6}$.
@@
Consider the convex geodesic (m1)(m-1)-gon made by $l_{1}$, $l_{2}$, $\dots$, $l_{m-1}$ that contains $\mathcal{P}_{m}$. We denote this polygon by $\mathcal{P}_{m-1}$, and let $\mathcal{T}$ be the triangle $\mathcal{P}_{m-1}\setminus\mathcal{P}_{m}$. In the previous example of $\mathcal{P}_{6}$, $\mathcal{T}$ is the green triangle in Figure 10. Let $\Theta_{3}$ be the sum of the $3$ angles for $\mathcal{T}$ as we defined previously, and let $\Theta_{m-1}$ be the sum of the m1m-1 angles for $\mathcal{P}_{m-1}$. In the first picture of Figure 10, $\Theta_{3}$ is the sum of the three angles marked in green. In the second picture, $\Theta_{m-1}$ is the sum of the five angles marked in blue.
@@
Now we add up the $m$ angles for $\mathcal{P}_{m}$ (whose sum equals $\Theta_{m}$) and the $3$ angles for $\mathcal{T}$ (whose sum is $2\pi-\sum\limits_{n^{+}\text{ or }n^{-}\in\mathcal{T}}\delta_{n}$ by Equation 2). Among these angles, there are two pairs of complementary angles, so their sum equals $2\pi$. These four angles appear in the two parallelograms dual to the two vertices that belong to both $\mathcal{P}_{m}$ and $\mathcal{T}$. In Figure 10, we draw these two parallelograms and mark the four angles in the third picture. Apart from these four angles, the sum of the remaining angles equals $\Theta_{m-1}$. Therefore, we obtain $\Theta_{m}+\Theta_{3}=\Theta_{m-1}+2\pi$.
@@
4 Polyhedra with $8$ Vertices
In this section, we focus on $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$, the space of centrally symmetric convex polyhedral surface with prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\delta_{3}$, and $\delta_{4}$. The goal is to show that every polyhedral surface in this space has a parallelogram decomposition that is invariant under the antipodal map. Restricted to this section, a loop arrangement always contains $6$ loops labeled by $l_{a}$, $l_{b}$, $l_{c}$, $l_{d}$, $l_{e}$, and $l_{f}$ on a sphere with $8$ vertices labeled by $\pm 1$, $\pm 2$, $\pm 3$ and $\pm 4$.

### 30 Mar 2026 v2

@@ Abstract
Let $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the moduli space of centrally symmetric convex polyhedral surfaces with $2N$ labeled vertices and prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$. We show that $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ has the structure of a real hyperbolic manifold of dimension 2N32N-3. When $N=4$ and 55, we show that every surface in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ can be decomposed into at most $2\binom{2N-2}{2}$ parallelograms, and the decomposition is invariant under the antipodal map. Using the edge-lengths of these parallelograms as coordinates, we show that the moduli space of centrally symmetric polyhedral surfaces with 88 unlabeled vertices and cone-deficits $\frac{\pi}{2}$ is isometric to the quotient of a real hyperbolic regular ideal 55-simplex by the dihedral group $D_{6}$.
@@ 1 Introducton
A polyhedron is convex if $0<\delta_{i}<2\pi$ for all ii. In [10], Thurston considered the moduli space of convex polyhedra with prescribed cone-deficits. By decomposing the surfaces of polyhedra into triangles, he built coordinates to describe these polyhedra, and showed that their moduli space admits a Hermitian metric, with respect to which it is a complex hyperbolic manifold. In addition, under certain conditions, the metric completion of this moduli space is an orbifold. This result provides a geometric way to interpret the monodromy groups of hypergeometric functions investigated by Picard [7][8], Mostow and Deligne [2][3][6]. In addition, this result has been applied in counting and enumerating specific structures on the sphere, such as counting convex tilings [5] and enumerating fullerenes [4]. Thurston’s idea in building coordinates and the Hermitian form is also used to study the moduli space of convex polygons [1].
@@
Inspired by this work, we study the space of centrally symmetric convex polyhedra. A polyhedron is centrally symmetric if it is isometric to a polyhedron symmetric about the origin, which means that a point $x$ lies on its surface if and only if its antipodal point x-x does. The map that sends $x$ to x-x for all $x$ on the surface is called the antipodal map. Thus, a centrally symmetric polyhedron has an even number $2N$ of vertices, and its cone-deficits are described by NN positive numbers that sum up to $2\pi$.
@@
The main results of this paper can be summarized as follows. In Section 2, we show that the moduli space of centrally symmetric convex polyhedra with prescribed cone-deficits has the structure of a real hyperbolic manifold. The way we compute the signature of the quadratic form is very similar to the way Thurston computed it for the Hermitian form in [10]. In addition, it is known that every centrally symmetric convex polygon can be decomposed into finitely many parallelograms. This leads us to conjecture that a similar result holds for every centrally symmetric convex polyhedral surface. In a previous work [12], we showed that the surface of every centrally symmetric octahedron has a parallelogram decomposition that is invariant under the antipodal map. In Section 3, we establish a method that relates a parallelogram decomposition of a polyhedral surface to a structure called “loop arrangement” on the sphere. Using this method, we generalize this result to centrally symmetric convex polyhedra with 88 and $10$ vertices in Section 4 and 5. In these cases, it turns out that the edge-lengths of the parallelograms in the decomposition can be used as coordinates to describe polyhedra. These coordinates are different from Thurston’s coordinates in [10], but may sometimes provide a more concrete geometric description of the moduli spaces when the polyhedra are centrally symmetric. For example, in Section 4, we show that the moduli space of centrally symmetric polyhedra with 88 vertices whose cone-deficits are $\frac{\pi}{2}$ is isometric to the quotient of a real hyperbolic regular ideal 55-simplex by a group isomorphic to $D_{6}$.
@@ 2 Space of Centrally Symmetric Convex Polyhedra
Let $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$ be NN positive numbers that sum up to $2\pi$. Let $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the space of centrally symmetric convex polyhedral surfaces with NN pairs of antipodal vertices labeled by $i^{+}$ and $i^{-}$ (where $1\leq i\leq N$) with prescribed cone-deficit $\delta_{i}$. In this space, two surfaces are considered equivalent if and only if they differ by a Euclidean isometry that preserves the labels of all the vertices.
@@
Next, let $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the moduli space of $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$, in which two polyhedral surfaces are equivalent if and only if they differ by a similarity preserving the vertex-labels. Alternatively, we can regard $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ as the space of polyhedral surfaces in $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ whose surface areas are equal to 11.
In [10], it was shown that the moduli space of convex polyhedra with NN prescribed cone-deficits has the structure of a complex hyperbolic manifold of dimension N3N-3. The main idea of proof is to show that the surface area is a Hermitian form of signature (1,n3)(1,n-3) with respect to the coordinates arising from decomposing the surfaces into triangles. The argument is based on induction on the number of vertices, where multiple vertices of a surface may “collide” to produce a surface with fewer vertices. One may refer to [10] and [9] for more details of this operation.
@@
Therefore, surfaces near $\mathscr{S}$ in $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ with area 11 correspond to vectors in $\mathbb{R}^{2N-2}$ of length 11 with respect to a quadratic form of signature (1,2N3)(1,2N-3). This establishes local coordinate charts from $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ to the real hyperbolic space $\mathbb{H}^{2N-3}$. Finally, any change-of-coordinate map on $\mathcal{C}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ keeps the surface area invariant. This shows that $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ is a real hyperbolic manifold of dimension 2N32N-3. ∎
Let $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ be the metric completion of $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ with respect to the real hyperbolic metric above. For example, the surface in the third picture of Figure 4 is not in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ but in $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ if it has area 11. This is because one can construct a Cauchy sequence of surfaces in $\mathcal{M}_{2N}(\delta_{1},\delta_{2},\dots,\delta_{N})$ converging to it by colliding two vertices and normalizing the surface area at the same time. We will return to this space in Section 4.
@@ 3 Loop Arrangements
Suppose that we are given $2N$ distinct vertices on $S^{2}$ labeled by $^{+}, 11^{-}$, $^{+}, 22^{-}$, $\dots$, $N^{+}$ and $N^{-}$, where $i^{+}$ and $i^{-}$ are antipodal for $1\leq i\leq N$. By a loop we mean a great circle on $S^{2}$ that does not pass through any labeled vertices. A set of loops form a loop arrangement if any two of them are non-homotopic with respect to the labeled vertices. Since loops are great circles, every two of them intersect exactly twice at two antipodal points. We assume without loss of generality that no three loops are concurrent, otherwise we can perturb any one of the them. As an example, in Figure 5, we sketch the “front view” of a loop arrangement with 66 loops on a sphere with 88 labeled vertices. Each red curve represents a loop. We use black dots to represent the vertices $1^{+}$, $2^{+}$, $3^{+}$, and $4^{+}$. In addition, we use gray dots to represent the vertices $2^{-}$, $3^{-}$, and $4^{-}$ (even though they should be invisible from the actual front view) because it will be helpful for our demonstrations later.
Figure 5: A loop arrangement with 66 loops (red) on a sphere with 88 labeled vertices.
@@
We first label all loops in $\mathscr{A}$. For example, if $\mathscr{A}$ is the loop arrangement in Figure 5, we can denote its 88 loops by $l_{a}$, $l_{b}$, $l_{c}$, $l_{d}$, $l_{e}$, and $l_{f}$. For convenience, these loops are marked by $a$, $b$, $\dots$, $f$ in the first picture of Figure 6. In the second picture, we draw the dual graph $\mathscr{D}^{*}$ of $\mathscr{A}$ in blue.
@@
Next, we assign angles to each quadrilateral. Let $Q$ be a quadrilateral in $\mathscr{D}^{*}$. Consider two edges of $Q$ that meet at a vertex $v$ in $Q$. We are going to assign a value $\theta$ as the angle made by these two edges in $Q$. To compute $\theta$, let $l_{i}$ and $l_{j}$ be the two loops to which these two edges are transverse. Then $l_{i}$ and $l_{j}$ divide the sphere into 44 lunes, one of which contains $v$ in its interior. We denote this lune by $\mathcal{L}$. For example, in the first picture of Figure 7, we sketch a quadrilateral $Q$ (blue), the angle $\theta$ we want to compute in $Q$, and the lune $\mathcal{L}$ (red) we consider to compute $\theta$.
@@
Since $2N-2\geq 4$, we can choose three distinct loops in $\mathscr{A}$, denoted by $l_{i}$, $l_{j}$ and $l_{k}$. By assumption, they are not concurrent, so they divide the sphere into 88 geodesic triangles. Denote one of these geodesic triangles by $\mathcal{T}$. Let $u_{i}$, $u_{j}$, and $u_{k}$ be the vertices of $\mathcal{T}$ such that $u_{i}$ is opposite to the edge of $\mathcal{T}$ contained in $l_{i}$, and similarly for $u_{j}$ and $u_{k}$. The first picture of Figure 8 is a sketch of $\mathcal{T}$ with its vertices.
@@
Now we focus on all the loops in $\mathscr{A}$. These loops divide $S^{2}$ into convex geodesic polygons. Let $\mathcal{P}_{m}$ be one of these polygons. Without loss of generality, we assume that it is bounded by the loops $l_{1}$, $l_{2}$, $\dots$, and $l_{m}$ in $\mathscr{A}$. Figure 9 shows an example when $m=6$. Consider the $m$ parallelograms in $\mathscr{D}^{*}$ dual to the $m$ vertices of $\mathcal{P}_{m}$. Let $v^{*}_{1}$, $v^{*}_{2}$, $\dots$, $v^{*}_{m}$ be the vertices of these parallelograms inside $\mathcal{P}_{m}$, which may be distinct or repeat. Let $\Theta_{m}$ be the sum of the $m$ angles of these parallelograms at the $v^{*}_{i}$’s. In the example in Figure 9, the 66 parallelograms are drawn in blue, and the sum of all the angles marked in red equals $\Theta_{6}$.
@@
Consider the convex geodesic (m1)(m-1)-gon made by $l_{1}$, $l_{2}$, $\dots$, $l_{m-1}$ that contains $\mathcal{P}_{m}$. We denote this polygon by $\mathcal{P}_{m-1}$, and let $\mathcal{T}$ be the triangle $\mathcal{P}_{m-1}\setminus\mathcal{P}_{m}$. In the previous example of $\mathcal{P}_{6}$, $\mathcal{T}$ is the green triangle in Figure 10. Let $\Theta_{3}$ be the sum of the 33 angles for $\mathcal{T}$ as we defined previously, and let $\Theta_{m-1}$ be the sum of the m1m-1 angles for $\mathcal{P}_{m-1}$. In the first picture of Figure 10, $\Theta_{3}$ is the sum of the three angles marked in green. In the second picture, $\Theta_{m-1}$ is the sum of the five angles marked in blue.
@@
Now we add up the $m$ angles for $\mathcal{P}_{m}$ (whose sum equals $\Theta_{m}$) and the 33 angles for $\mathcal{T}$ (whose sum is $2\pi-\sum\limits_{n^{+}\text{ or }n^{-}\in\mathcal{T}}\delta_{n}$ by Equation 2). Among these angles, there are two pairs of complementary angles, so their sum equals $2\pi$. These four angles appear in the two parallelograms dual to the two vertices that belong to both $\mathcal{P}_{m}$ and $\mathcal{T}$. In Figure 10, we draw these two parallelograms and mark the four angles in the third picture. Apart from these four angles, the sum of the remaining angles equals $\Theta_{m-1}$. Therefore, we obtain $\Theta_{m}+\Theta_{3}=\Theta_{m-1}+2\pi$.
@@
4 Polyhedra with 88 Vertices
In this section, we focus on $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$, the space of centrally symmetric convex polyhedral surface with prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\delta_{3}$, and $\delta_{4}$. The goal is to show that every polyhedral surface in this space has a parallelogram decomposition that is invariant under the antipodal map. Restricted to this section, a loop arrangement always contains 66 loops labeled by $l_{a}$, $l_{b}$, $l_{c}$, $l_{d}$, $l_{e}$, and $l_{f}$ on a sphere with 88 vertices labeled by $\pm 1$, $\pm 2$, $\pm 3$ and $\pm 4$.